OpenStudy (hannahwaller5):

Will Give Medal! Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately how long does it take until the soccer ball hits the ground again? 1.3 secs 2.6 secs 16.0 secs 27.6 secs

OpenStudy (yumyum247):

Is this a physics question???

OpenStudy (hannahwaller5):

No its Algebra 1

OpenStudy (anthonyym):

I would say it is because it's kinematics

OpenStudy (anonymous):

at the highest point velocity v=0 g=-32 ft/sec^2 initial velocity u=42 ft/sec v=u+gt 0=42-32t 32 t=42 \[t=\frac{ 42 }{ 32 }=\frac{ 21 }{ 16 }\] total time of flight=2t \[=2 \times \frac{ 21 }{ 16 }=\frac{ 21 }{ 8 } \sec=?\]

OpenStudy (yumyum247):

U guessed it right...

OpenStudy (hannahwaller5):

So it would be b... because 21/8 is 2.625

OpenStudy (anonymous):

correct.

OpenStudy (hannahwaller5):

Thanks So Much!!! (:

OpenStudy (anonymous):

yw